calculate the directional derivative of scalar field f(x,y,z)=(x ²+yz+z²)at (1,2,1) along n^=1/3(2i-j+k^)

1
Expert's answer
2022-01-07T09:29:46-0500

The directional derivative in the direction of "\mathbf{n} = \dfrac23\mathbf{i}-\dfrac13\mathbf{j} + \dfrac13\mathbf{k}" is given as follows:


"\nabla_{\mathbf{n}}f = \nabla f\cdot \dfrac{\mathbf{n}}{|\mathbf{n}|}"

where "\nabla f" is the gradient of "f". Thus, obtain:


"\nabla f = 2x\mathbf{i} + z\mathbf{j}+(y+2z)\mathbf{k}\\
|\mathbf{n}| = \sqrt{\dfrac49 + \dfrac19 + \dfrac19} = \dfrac{\sqrt{6}}{3}\\
\dfrac{\mathbf{n}}{|\mathbf{n}|} = \dfrac{1}{\sqrt{6}}(2\mathbf{i}-\mathbf{j} + \mathbf{k})"

Finally,


"\nabla_{\mathbf{n}}f = (2x\mathbf{i} + z\mathbf{j}+(y+2z)\mathbf{k})\cdot \dfrac{1}{\sqrt{6}}(2\mathbf{i}-\mathbf{j} + \mathbf{k})=\\
=\dfrac{1}{\sqrt{6}}(4x-z+y+2z) = \dfrac{1}{\sqrt{6}}(4x+y+z)"

Substituting point (1,2,1), obtain:


"\dfrac{1}{\sqrt{6}}(4\cdot 1+2+1) = \dfrac{7}{\sqrt{6}}"

Answer. "\dfrac{7}{\sqrt{6}}".